John is 6 years older than his brother. He will be twice as old as his brother in 4 years. How old is John now?
Answer by macston(5194)

B=brother's age;J=John's age=B+6
J+4=2(B+4) Substitute for J
(B+6)+4=2(B+4)
B+10=2B+8 Subtract 8 from each side
B+2=2B Subtract B from each side
2=B John's brother is 2 years old
J=B+6=8 ANSWER John is 8 years old
CHECK
J+4=2(B+4)
8+4=2(2+4)
12=12